# Enhanced superconductivity in LaRu3Si2 by chemical pressure tuning of kagome flat bands

### v1

@@ Abstract
We measure specific heat by the relaxation technique in a Quantum Design Physical Properties Measurement System (PPMS). For each measurement, the addenda (with Apiezon N grease mounted on the stage) are measured before setting the sample for the main measurement. To eliminate contributions from non-magnetic impurity Ru metal, we subtract $\gamma_{\mathrm{Ru}}=2.8\,\mathrm{mJ/molK^{2}}$ and $\beta_{\mathrm{Ru}}=0.068\,\mathrm{mJ/molK^{4}}$ from the total $\gamma$ and $\beta$ values Li et al. (2011); Reese and Johnson (1970). For each value, we multiply the ratio of Ru impurity $y/3$, where $y$ represents the ratio of excess Ru, namely $y=1.5$ for $x=0,0.02,0.05$ and $y=2$ for $x=0.07$. We summarize $\gamma$ and $\beta$ values before and after the subtraction of the Ru contribution in Table 1. Errors for the fitting parameters $\beta$ and $\gamma$ in Fig. 3(d) are estimated by the bootstrap method. We estimate the density of states (DOS) at the Fermi energy, $N(E_{\mathrm{F}})$, using the relation
@@
which has units of $\mathrm{J}^{-1}$. We convert this to $\mathrm{eV}^{-1}$ per unit cell (u.c.). Furthermore, since $\gamma$ is calculated assuming a spinless electronic structure, we divide by 2 to obtain $N(E_{\mathrm{F}})$ in units of $\mathrm{eV}^{-1}$ per unit cell per spin Tinkham (2004); Kriener et al. (2022).
For the superconducting state, the specific heat can be fitted by an analytical expression within the framework of the BCS theory. Here we approximate the temperature dependence to be linear below $T_{\mathrm{c}}$, which is valid for our temperature range ($>2\,$K) Gui and Cava (2022). This method allows us to precisely determine $T_{\mathrm{c}}$ by the equal-area construction, which is crucial for tracking the evolution of $T_{\mathrm{c}}$ in our crystals: We first numerically calculate the area under the curve of $\Delta C/T=C/T-C_{\mathrm{N}}/T$ below 888\,8K, denoted as $S$ [Fig. S3(a)]. We then fit $\Delta C/T$ below 666\,6K by a linear function $\Delta C/T=aT+b$. Imposing entropy conservation, the area under the curve is thus given by
@@
where $T_{0}=-b/a$ is the intercept of the linear fit, and $T_{\mathrm{min}}$ is the lowest temperature at which data were taken. The first term corresponds to the area for $T_{0}\leq T\leq T_{\mathrm{c}}$ and the second to that for $T_{\mathrm{min}}\leq T\leq T_{\mathrm{c}}$, as illustrated in Fig. S3(b). Solving this equation, we analytically determine the transition temperature as
@@
We provide the calculated ratio $\Delta C(T_{\mathrm{c}})/\gamma T_{\mathrm{c}}$ in Table 1. A slight deviation from the BCS value of $1.43$ may suggest a moderate superconducting coupling strength. Given that the Sommerfeld coefficient of pure Ru is small ($\gamma_{\mathrm{Ru}}=2.8\,\mathrm{mJ/mol\,K^{2}}$), the residual specific heat contribution from the Ru impurity is expected to be negligible.
To explore the effects of $\beta$ and $\gamma$ on $T_{\mathrm{c}}$, we consider the McMillan equation
$T_{\mathrm{c}}=\frac{\Theta_{\mathrm{D}}}{1.45}\exp\left(\frac{-1.04(1+\lambda)}{\lambda-\mu^{*}(1+0.62\lambda)}\right),$ (S5)
where $\Theta_{\mathrm{D}}$ is the Debye temperature, $\lambda$ is the electron-phonon coupling constant, and $\mu^{*}$ is the Coulomb pseudopotential Carbotte (1990). The electron-phonon coupling $\lambda$ is expressed in terms of the density of states at the Fermi level, $N(E_{\mathrm{F}})$, as
$N(E_{\mathrm{F}})V$ $=\frac{\lambda-\mu^{*}}{1+\lambda}$ (S6)
$=1-\frac{1+\mu^{*}}{1+\lambda},$ (S7)
where $V$ denotes the pairing potential. This relation indicates that an increase in $\gamma$ will also enhance $\lambda$, thereby promoting $T_{\mathrm{c}}$; this is assuming a relatively static value of $V$. Meanwhile, the Debye temperature $\Theta_{\mathrm{D}}$ is related to $\beta$ viz.
$\Theta_{\mathrm{D}}=\left(\frac{12\pi^{4}Nk_{B}}{5\beta}\right)^{\frac{1}{3}},$ (S8)
where $N$ is the number of atoms per formula unit. Since $\Theta_{\mathrm{D}}$ scales inversely with $\beta^{1/3}$, an increase in $\beta$ suppresses $\Theta_{\mathrm{D}}$, leading to a reduction in $T_{\mathrm{c}}$. Overall, while both $\beta$ and $\gamma$ influence $T_{\mathrm{c}}$, the observed enhancement of $T_{\mathrm{c}}$ suggests that the effect of $\gamma$ dominates. Setting $\mu^{*}=0.13$, we estimate $\lambda$ to be $0.63\sim 0.66$ from Eq. (S5), as summarized in Table 1. The enhanced electron-phonon coupling constant $\lambda>0.5$ suggests a moderately coupled superconductor. This is consistent with $\Delta C(T_{\mathrm{c}})/\gamma T_{\mathrm{c}}>1.43$.
@@
Table 1: Properties of LaRu3(Si1-xGex)2 probed by specific heat measurements. For the electron and phonon contributions $\gamma$ and $\beta$, values before and after subtracting a subtle contribution from Ru impurity in LaRu3(Si1-xGex)2 + $y$Ru are given. $\Delta C(T_{\mathrm{c}})$ denotes the specific heat jump at $T_{\mathrm{c}}$, and $\Delta C(T_{\mathrm{c}})/\gamma T_{\mathrm{c}}=1.43$ holds for weakly coupled superconductors. $N(E_{\mathrm{F}})$ and $\lambda$ denote the DOS and the electron-phonon coupling constant, respectively.
@@
$\gamma-y/3\cdot\gamma_{\mathrm{Ru}}\,\,\mathrm{(mJ/molK^{2})}$ 36.9 36.0 39.3 42.0
@@
$\beta-y/3\cdot\beta_{\mathrm{Ru}}\,\,\mathrm{(mJ/molK^{4})}$ 0.18 0.15 0.21 0.20
$\Delta C(T_{\mathrm{c}})/\gamma T_{c}$ 1.42 1.64 1.54 1.57

### v2

@@ Abstract
We measure specific heat by the relaxation technique in a Quantum Design Physical Properties Measurement System (PPMS). For each measurement, the addenda (with Apiezon N grease mounted on the stage) are measured before setting the sample for the main measurement. To eliminate contributions from non-magnetic impurity Ru metal, we subtract $\gamma_{\mathrm{Ru}}=2.8\,\mathrm{mJ/molK^{2}}$ and $\beta_{\mathrm{Ru}}=0.068\,\mathrm{mJ/molK^{4}}$ from the total $\gamma$ and $\beta$ values Li et al. (2011); Reese and Johnson (1970). For each value, we multiply the ratio of Ru impurity $y/3$, where $y$ represents the ratio of excess Ru, namely $y=1.5$ for $x=0,0.02,0.05$ and $y=2$ for $x=0.07$. We summarize $\gamma$ and $\beta$ values before and after the subtraction of the Ru contribution in Table S1.
Errors for the fitting parameters $\beta$ and $\gamma$ are estimated by the bootstrap method. This allows us to evaluate errors or the stability of fitting while incorporating systematic measurement errors. The procedure is as follows: We first fit the data and calculate residuals. We then compute the total measurement error for each data point by performing the relaxation method measurement 3 times at each temperature and then calculating the mean and standard deviation $\sigma(T)$. Next, we randomly sample values from zero to $\sigma(T)$, add them to the experimental data and obtain fit parameters for this perturbed data. This is repeated 100 times. We take an average of these trials as a central value for $\beta$ and $\gamma$; we also use its standard deviation as a fit error for these parameters. Other errors, such as sample mass and batch dependence, are ignored as we precisely measured the mass and uniformly melted all samples.
S VI Electronic and phononic specific heat
We estimate the density of states (DOS) at the Fermi energy, $N(E_{\mathrm{F}})$, using the relation
@@
which has units of $\mathrm{J}^{-1}$. We convert this to $\mathrm{eV}^{-1}$ per unit cell (u.c.). Furthermore, since $\gamma$ is calculated assuming a spinless electronic structure, we divide by 2 to obtain $N(E_{\mathrm{F}})$ in units of $\mathrm{eV}^{-1}$ per unit cell per spin Tinkham (2004); Kriener et al. (2022). The errors of $N(E_{\mathrm{F}})$ in Fig. 3 (e) are estimated by the error propagation of $\gamma$. In Table  S1, we also present the Debye temperature $\Theta_{\mathrm{D}}$ calculated from $\beta$ as
$\Theta_{\mathrm{D}}=\left(\frac{12\pi^{4}Nk_{\mathrm{B}}}{5\beta}\right).$ (S2)
In the Main Text, we discuss the observed reduction of $\Theta_{D}$ which is summarized in Table S1. Assuming Eq. (3), of the Main Text, this may be understood to imply a softening of $\omega\right>$ with $x$ for the phonon modes relevant to the superconducting $T_{\mathrm{c}}$. However, softening appears unreasonable considering the dominant role of the $B_{3u}$ phonon in DFT Deng et al. (2025), its nature as a vibration in the hexagonal $ab$ plane, and the lack of a lattice expansion along $a$ in Fig. 2(c). The observed softening of $\Theta_{\mathrm{D}}$ with $x$ should then be attributed to the elastic coupling between kagome layers, which is irrelevant to the superconductivity Deng et al. (2025).
S VII Specific heat jump at $T_{\mathrm{c}}$
For the superconducting state, the specific heat can be fitted by an analytical expression within the framework of the BCS theory. Here, we approximate the temperature dependence to be linear below $T_{\mathrm{c}}$, which is valid for our temperature range ($>2\,$K) Gui and Cava (2022). This method allows us to precisely determine $T_{\mathrm{c}}$ by the equal-area construction, which is crucial for tracking the evolution of $T_{\mathrm{c}}$ in our crystals: We first numerically calculate the area under the curve of $\Delta C/T=C/T-C_{\mathrm{N}}/T$ below 888\,8K, denoted as $S$ [Fig. S4(a)]. We then fit $\Delta C/T$ below 666\,6K by a linear function $\Delta C/T=aT+b$. Imposing entropy conservation, the area under the curve is thus given by
@@
where $T_{0}=-b/a$ is the intercept of the linear fit, and $T_{\mathrm{min}}$ is the lowest temperature at which data are taken. The first term corresponds to the area for $T_{0}\leq T\leq T_{\mathrm{c}}$ and the second to that for $T_{\mathrm{min}}\leq T\leq T_{\mathrm{c}}$, as illustrated in Fig. S4(b). Solving this equation, we analytically determine the transition temperature as
@@
The errors of $a$ and $b$ are evaluated by the same method as for $\beta$, $\gamma$, and $T_{\mathrm{c}}$ described above and shown in Fig. 3. Error propagation was carefully considered.
We provide the calculated ratio $\Delta C(T_{\mathrm{c}})/\gamma T_{\mathrm{c}}$ in Table S1. A slight deviation from the BCS value of $1.43$ may suggest a moderate superconducting coupling strength. This behavior is consistent with a previous report Chakrabortty et al. (2023). Given that the Sommerfeld coefficient of pure Ru is small ($\gamma_{\mathrm{Ru}}=2.8\,\mathrm{mJ/mol\,K^{2}}$), the residual specific heat contribution from the Ru impurity is expected to be negligible.
S VIII Electron-phonon coupling strength and bare density of states
Using Eq. (1) and Eq. (3) of the Main Text, we estimate the electron-phonon coupling constant $\lambda$. Setting $\mu^{*}=0.13$, we estimate $\lambda$ to be $0.63\sim 0.66$, as summarized in Table S1. The enhanced electron-phonon coupling constant $\lambda>0.5$ suggests a moderately coupled superconductor. This is consistent with $\Delta C(T_{\mathrm{c}})/\gamma T_{\mathrm{c}}>1.43$ in Fig. 4(h).
@@
which allows for comparison to band structure calculations. Here, $N_{\mathrm{bare}}(E_{\mathrm{F}})$ has the same unit as $N(E_{\mathrm{F}})$ because $\lambda$ is dimensionless. This yields approximately $N_{\mathrm{bare}}(E_{\mathrm{F}})=5\,\mathrm{/eV/u.c/spin}$ and reasonably corresponds to the calculated DOS at $E=-0.1\,\mathrm{eV}$ in Fig. 3(f).
S IX Strength of doping effect on $T_{\mathrm{c}}$
@@
Table S1: Properties of LaRu3(Si1-xGex)2 probed by specific heat measurements. For the electron and phonon contributions $\gamma$ and $\beta$, values before and after subtracting a subtle contribution from Ru impurity in LaRu3(Si1-xGex)2 + $y$Ru are given. $\Delta C(T_{\mathrm{c}})$ denotes the specific heat jump at $T_{\mathrm{c}}$, and $\Delta C(T_{\mathrm{c}})/\gamma T_{\mathrm{c}}=1.43$ holds for weakly coupled superconductors. $N(E_{\mathrm{F}})$ and $\lambda$ denote the DOS and the electron-phonon coupling constant, respectively.
@@
$\gamma^{\prime}=\gamma-y/3\cdot\gamma_{\mathrm{Ru}}\,\,\mathrm{(mJ/molK^{2})}$ 36.9 36.0 39.3 42.0
@@
$\beta^{\prime}=\beta-y/3\cdot\beta_{\mathrm{Ru}}\,\,\mathrm{(mJ/molK^{4})}$ 0.18 0.15 0.21 0.20
$\Theta_{\mathrm{D}}$ (K) 405(6) 423(20) 379(6) 384(9)
$\Delta C(T_{\mathrm{c}})/\gamma^{\prime}T_{c}$ 1.42 1.64 1.54 1.57
@@
Material $T_{\mathrm{c}}(0)\,$(K) $x_{0}$(%) $T_{\mathrm{c}}(x_{0})$(K) $\Delta T_{\mathrm{c}}$(%) $\Delta T_{\mathrm{c}}/x_{0}$ $N(E_{\mathrm{F}})_{\mathrm{exp}}$ $N(E_{\mathrm{F}})_{\mathrm{DFT}}$ reference
InTe1-xSex $2.6$ $30$ $5.0$ $92$ $3$$\times$ $\times$  Kriener et al. (2018)
1T-Pd(Te1-xSex) $1.64$ $50$ $2.7$ $67$ $1$bigcirc $\times$  Liu et al. (2021)
Sn1-xGexTe $2.1$ $1.2$ $1.3$ $-38$ $-32$$\times$ $\times$  Grassie and Benyon (1972)
Sn1-xPbxTe $2.1$ $12$ $1.9$ $-10$ $-0.8$$\times$ $\times$  Grassie and Benyon (1972)
K0.8Fe2(Se1-xTex)2 $31$ $25$ $14.9$ $-52$ $-2$$\times$ $\times$  Ozaki et al. (2013)
[…]
Cs(V1-xNbx)3Sb5 $2.9$ $7$ $4.5$ $55$ $8$$\times$ $\times$  Li et al. (2022)
Cs(V1-xTax)3Sb5 $3.1$ $16$ $5.3$ $71$ $4$$\times$ $\times$  Liu et al. (2024b)
[…]